OpenStudy (anonymous):

\[\int\limits_{1}^{4}{1 \over \sqrt{x}+x}dx\]

OpenStudy (bahrom7893):

multiply top and bottom by sqrt(x)-x i think..

OpenStudy (ash2326):

let x be tan^2 \(\theta\) dx= 2 tan \(\theta\) sec^2 \(\theta\) when x=1 tan^2 \(\theta\)= 1 \(\theta\)= \(\pi/4\) when x=4 tan^2 \(\theta\)=4 \(\theta\)= \(\tan^{-1} 2\) so we have the integral now \[\int_{\theta=\pi/4}^{\theta=\ tan^{-1} 2} \frac{ 2 tan (\theta ) sec^2 (\theta)} { tan \theta+ tan^2 \theta } d\theta \] now let's cancel the tan \(\theta\) from numerator and denominator we get \[\int_{\theta=\pi/4}^{\theta=\ tan^{-1} 2} \frac{ 2 sec^2 (\theta)} { tan \theta+ 1} d\theta \] let's substitute ( 1+ tan \(\theta\)) as t so we get sec^2 \(\theta\) d\(\theta\)= dt now when \(\theta\)= \(\pi/4\) t= 2 and when \(\theta\)= \(\tan^{-1} 2\) t= 3 so now we have integral as \[ \int_{t=2} ^{t=3} \frac{2}{t} dt\] we know that integral of 1/t is log t (with base e) so we get \[ [2 \log_{e} t]_{t=2}^{t=3} \] we get finally \[ 2 \log_{e} 3-2 \log_{e} 2\] or \[2\log_{e} {\frac{3}{2}}\]

OpenStudy (nenadmatematika):